OpenStudy (anonymous):

A pendulum oscillates in a vertical plane, length of string is 1m. At what point vertically below the point of suspension must a nail be fixed so that bob starts revoluting about the nail as its centre? Pendulum makes 60degree with the vertical. Pl.answer with explanation

OpenStudy (anonymous):

are u asking for the point where there is maximum oscilaton of the pendulum when the nail is fixed?

OpenStudy (anonymous):

 initial potential energy of bob = mg(1/2) = kinetic energy at lowest point = (1/2)mv^2 frm here v = \[\sqrt{10}\] now to complete a circle a speed to \[\sqrt{5gx}\] (radius of circle) is required.. \[\sqrt{5gx}\] = \[\sqrt{10}\] => x = 0.2 (from lowest point).